OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

OpenStudy (anonymous):

13x - 6y = 20.............i 7x + 4y = 18.............ii

OpenStudy (anonymous):

Now, (20+6y)/13=(18-4y)/7

OpenStudy (anonymous):

Now, substitute all the values and see which one satisfies

OpenStudy (anonymous):

Does it help?

OpenStudy (moonlitfate):

Do you still need help with this problem? :)

OpenStudy (anonymous):

Isolate both x's and use x1= x2 to solve for y

OpenStudy (anonymous):

i got y = -2 - .6 :/

OpenStudy (anonymous):

you guys did it a very complicated way^ i still don't understand..

jimthompson5910 (jim_thompson5910):

Multiply both sides of 13x - 6y = 20 by 2 to get 13x - 6y = 20 2(13x - 6y) = 2*20 26x - 12y = 40 Multiply both sides of 7x + 4y = 18 by 3 to get 7x + 4y = 18 3(7x + 4y) = 3*18 21x + 12y = 54 So we now have the system 26x - 12y = 40 21x + 12y = 54

jimthompson5910 (jim_thompson5910):

Add the equations 26x - 12y = 40 21x + 12y = 54 --------------- 47x = 94 and solve for x 47x = 94 x = 94/47 x = 2

OpenStudy (anonymous):

Haha, I was too lazy to write all that.

jimthompson5910 (jim_thompson5910):

then use this to find y 13x - 6y = 20 13(2) - 6y = 20 26 - 6y = 20 -6y = 20 - 26 -6y = -6 y = -6/(-6) y = 1

OpenStudy (anonymous):

Oh i understand better now, why did you multiply both sides by 2 in the first step? how did you get that number. i didn't know you had to add the equations together either.

OpenStudy (anonymous):

lol c=

OpenStudy (precal):

that is called elimination. we use this method to eliminate either x or y. Since you wanted the y value of the system, he eliminated your x

OpenStudy (precal):

I will post a couple of examples for you

OpenStudy (precal):